EXACT ENUMERATION OF 1342-AVOIDING PERMUTATIONS: A CLOSE LINK WITH LABELED TREES AND PLANAR MAPS

MIKLÓS BÓNA

Abstract. Solving the first nonmonotonic, longer-than-three instance of a classic enumeration problem, we obtain the generating function H(x) of all 1342-avoiding permutations of length n as well as an exact formula for their number S_{n}(1342). While achieving this, we bijectively prove that the number of indecomposable 1342avoiding permutations of length n equals that of labeled plane trees of a certain type on n vertices recently enumerated by Cori, Jacquard and Schaeffer, which is in turn known to be equal to the number of rooted bicubic maps enumerated by Tutte in 1963. Moreover, H(x) turns out to be algebraic, proving the first nonmonotonic, longer-than-three instance of a conjecture of Zeilberger and Noonan. We also prove that \sqrt[n]{S_{n}(1342)} converges to 8, so in particular, lim_{n→∞}(S_{n}(1342)/S_{n}(1234)) = 0.